What is the percent yield of COβ‚‚(𝑔) if 46.5 grams of COβ‚‚(𝑔) is recovered from the reaction of 21.5 grams of Cβ‚‚Hβ‚‚(𝑔) according to the reaction below?Β 


2C2H2(g)+5 O2(g)β€”> 4CO2(g)+2H2O(g)


1
Expert's answer
2021-04-15T02:03:02-0400

Moles of Cβ‚‚Hβ‚‚(𝑔)Β = "\frac{21.5}{26}=0.83 moles"



Moles of COβ‚‚(𝑔)Β = "\frac{4×0.83}{2}=1.66moles"



Mass of CO2= "1.66\u00d7 44=73.04moles"



Percent yield = "\frac{46.5}{73.04}×100=63.67" %